OpenStudy (anonymous):

Solve y over 2− 3 = 12 a:30 b;−30 c;18 d;−18

OpenStudy (anonymous):

\[\frac{ y }{ 2 - 3 } = \frac{ y }{ -1 }\] and \[\frac{ y }{ -1 } = 12\] multiply each side by -1 \[y = 12\]

OpenStudy (anonymous):

you sure you copied that right?

OpenStudy (anonymous):

or did you mean \[\frac{ y }{ 2 } - 3 = 12\]